OpenStudy (anonymous):

differentiate (t-1/t)(t^2-1/t^2)

OpenStudy (anonymous):

no look more closely

OpenStudy (anonymous):

Is it \((t-\frac{1}{t})(t^2-\frac{1}{t^2})\) in your question?

OpenStudy (anonymous):

thats better

OpenStudy (anonymous):

If it is the case, I suggest you to expand the factor first.

OpenStudy (anonymous):

if i knew how to complete that succesfully i wouldn't have put my question on open study

OpenStudy (anonymous):

Can you try expanding the factors first?

OpenStudy (anonymous):

You can also use product rule and power rule. \[(f\cdot g) ^\prime (x) = f^\prime g + f g^\prime\] \[f = x^n \Rightarrow f^\prime (x) =n \cdot x^{n-1} \]

OpenStudy (anonymous):

Hint: \((t-\frac{1}{t})(t^2-\frac{1}{t^2})\) = \(t(t^2-\frac{1}{t^2}) - \frac{1}{t}(t^2-\frac{1}{t^2})\)

OpenStudy (anonymous):

yes that i have done but i,m stuck just a little further up than that

OpenStudy (anonymous):

The reason why I'm not using product rule is because expanding it seems to be a easier way. But you can use the product rule of course.

OpenStudy (anonymous):

and so what next

OpenStudy (anonymous):

What have you got by expanding it?

OpenStudy (anonymous):

t^3-1/t^3-1/t^2+1/t^2

OpenStudy (anonymous):

make the last product 1/t^3

OpenStudy (anonymous):

both 1/t^3 cancel and i,m left with t^3-1/t^2

OpenStudy (anonymous):

You'd better check your calculation again..

OpenStudy (anonymous):

i have several times

OpenStudy (anonymous):

\[t(t^2-\frac{1}{t^2}) - \frac{1}{t}(t^2-\frac{1}{t^2})\]\[=t(t^2)-(t)(\frac{1}{t^2}) - \frac{1}{t}(t^2)-(\frac{1}{t})(\frac{1}{t^2})\]\[=...?\]

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

And can you simplify that?

OpenStudy (anonymous):

t^3-1/t^3-1/t^3-2/t^3

OpenStudy (anonymous):

how's the calculus coming along?

OpenStudy (anonymous):

\[t (\frac{1}{t^2}) = \frac{t}{t^2} = \frac{1}{t}\], not 1/t^3

OpenStudy (anonymous):

galad to see you algabraic

OpenStudy (anonymous):

so close.

OpenStudy (anonymous):

Wow, should I go now? Btw, (1/t) (t^2) =/= 1/t^3, (1/t)(t^2) = t

OpenStudy (anonymous):

no thats not the answer

OpenStudy (anonymous):

http://i.imgur.com/CMgMz.gif

OpenStudy (anonymous):

I'm not saying that it is the answer, I haven't started differentiation yet..

OpenStudy (anonymous):

\[(t-\frac{1}{t})(t^2-\frac{1}{t^2})\]\[=(t^3-t-\frac{1}{t}+\frac{1}{t^3})\] So, \[\frac{d}{dt}(t-\frac{1}{t})(t^2-\frac{1}{t^2})\]\[=\frac{d}{dt}(t^3-t-\frac{1}{t}+\frac{1}{t^3})\]Use power rules, you can do it.

OpenStudy (anonymous):

Sorry, it should be use the power rule :\

OpenStudy (anonymous):

once again so close

OpenStudy (anonymous):

You have to differentiate it yourself. That is not the answer.

OpenStudy (anonymous):

i'm having trouble with the second halve

OpenStudy (anonymous):

Which one?

OpenStudy (anonymous):

i have 3t^2-1.......?

OpenStudy (anonymous):

How did you get it?

OpenStudy (anonymous):

the power rule

OpenStudy (anonymous):

it is the answer but the rest is not adding up correctly

OpenStudy (anonymous):

Power rule: d\dx (x^n) = nx^(n-1)

OpenStudy (anonymous):

sorry make that simple differentiation

OpenStudy (anonymous):

x^2 = 2x derivative

OpenStudy (anonymous):

d\dx (x^2) = 2x => correct What about d/dx (x^3) ?

OpenStudy (anonymous):

x^3 = 3x^2

OpenStudy (anonymous):

Yup, so what about d/dt (t^3) ?

OpenStudy (anonymous):

my times up here i have to move on

OpenStudy (anonymous):

thanks for confusing me

OpenStudy (anonymous):

Sarcastic... When you get that, you can differentiate it term by term, it shouldn't be difficult when you use the power rule. It would be even more terrible to use product rule.